Cl2(g) ⇌ 2Cl(g)

What is happening to the entropy in the following reaction? If entropy changes, what is the sign of ΔS?

H2O(g) ⇌ ½ O2(g) + H2(g)

What is happening to the entropy in the following reaction? If entropy changes, what is the sign of ΔS?

Ag+(aq) + Cl-(aq) ⇌ AgCl(s)

What is happening to the entropy in the following reaction? If entropy changes, what is the sign of ΔS?
Answer #1

Cl2(g) ⇌ 2Cl(g)

\DeltaS   = S0f products - S0 f reactants

       = 2*165.198-223.066

        = 107.33J/K

\DeltaS is positive.

The number of partilces are increases randomness is increases. The entropy change is increases. \Delta S is +ve

H2O(g) ⇌ ½ O2(g) + H2(g)

\DeltaS   = S0f products - S0 f reactants

          = 130.684+1/2*205-188.825

           = 130.684+102.5-188.825

         = 44.359J/K

\DeltaS is positive.

The number of partilces are increases randomness is increases. The entropy change is increases. \Delta S is +ve

Ag+(aq) + Cl-(aq) ⇌ AgCl(s)

\DeltaS   = S0f products - S0 f reactants

       = 96.2-(72.676+56.484)

       = -32.96J/K

\DeltaS is -ve

The aqueous solution convert to solid randomness is decreases. The entropy change is decreases. \Delta S is -ve